A collection of 33 coins, consisting of nickels, dimes, and quarters, has a value of $3.30. If there are three times as many nickels as quarters, and one-half as many dimes as nickels, how many coins of each kind are there?
Answer by mananth(16949)

A collection of 33 coins, consisting of nickels, dimes, and quarters, has a value of $3.30. If there are three times as many nickels as quarters, and one-half as many dimes as nickels, how many coins of each kind are there?
...
quarters = x
nickels = 3x
dimes=3x/2
..
Their sum = 33
x+3x+3x/2 =33
11x/2=33
11x=66
x=6 quarters
nickels =18
dimes=9
CHECK
18*5+9*10+6*25
90+90+150=330
value=3.30
m.ananth@hotmail.ca